# UFID: A Unified Framework for Input-level Backdoor Detection on Diffusion Models

### v1

@@ Appendix F More Details about Score Distributions
$\begin{split}\Sigma_{\tilde{\theta}}(x^{\prime}_{t},t)&=1/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(\frac{\alpha_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}})=1/(\frac{\alpha_{t}-\bar{\alpha}_{t}+\beta_{t}}{\beta_{t}(1-\bar{\alpha}_{t-1})})=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}\\ \mu_{\tilde{\theta}}(x^{\prime}_{t},t)&=\frac{1}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}}(\frac{\sqrt{\alpha_{t}}}{\beta_{t}}x^{\prime}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}}{1-\bar{\alpha}_{t-1}}x_{0})/\frac{1}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}}(\frac{\alpha_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}})\\ &=(\frac{\sqrt{\alpha_{t}}}{\beta_{t}}x^{\prime}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}}{1-\bar{\alpha}_{t-1}}x_{0})\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}\\ &=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}x^{\prime}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}x_{0}\\ &=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}x^{\prime}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\frac{1}{\sqrt{\bar{\alpha}_{t}}}(x^{\prime}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}\boldsymbol{\epsilon}_{t})\\ &=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x^{\prime}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}\boldsymbol{\epsilon}_{t}\Big{)}\end{split}$ (11)
According to Lemma 3.2, if we add Gaussian noise to the origin input image, which results in $\mathcal{N}(0,\rho^{2}I)$, then the distribution of generated images of the diffusion model has the same mean, but a variance scaled by $\frac{1}{\rho^{2}}$, where $\rho^{2}=2$ for clean samples.
@@
Definition G.5 (Backdoor Forward process).
Let $x_{0}^{b}\sim q(x_{b})$ denote a sample from target data distribution, $\delta$ denote a trigger, and $x_{T}^{b}\sim\mathcal{N}(\delta,I)$ denote the pure Gaussian noise attached by a trigger. Given the variance schedule $\{\beta_{t}\}_{t=1}^{T}$ in DDPM [22], define the forward process to diffuse $x_{0}^{b}$ to $x_{T}^{b}$ for backdoor samples:
@@
For D2I attack, the target generated image is specific (e.g., Mickey Mouse), thus its variance is 0. For In-D2D and Out-D2D, the target images belong to the same class, hence their variance should be much lower than that of images generated from the clean model, as those are from various, completely different classes. Therefore, under the above case and Lemma 3.2, the distributions generated by clean and backdoor samples after noise addition are $\mathcal{N}(x_{c},\sigma_{c}\frac{1}{\rho^{2}}I)$ and $\mathcal{N}(x_{b},\sigma_{b}\frac{1}{\rho^{2}}I)$, respectively. In particular, we consider three attack methods. Hence, we have three kinds of $\sigma_{b}$. Let $\sigma_{D2I}$ denote the variance of D2I attack, $\sigma_{iD2D}$ denote the variance of In-D2I attack, $\sigma_{oD2I}$ denote the variance of Out-D2I attack defined in Appendix  B. We then have,
$0=\frac{1}{\rho^{2}}\sigma_{D2I}\ll\frac{1}{\rho^{2}}\sigma_{In-D2D},\frac{1}{\rho^{2}}\sigma_{Out-D2D}\ll\frac{1}{\rho^{2}}\sigma_{c}.$ (14)
Equation 14 shows the different variances of the images generated from the fixed diffusion model $f_{\theta}$ with noise added to the input in terms of clean and backdoor samples.
Under Lemma 3.1, which states that two points sample from a distribution of larger variance tend to have a larger distance than those from a distribution of lower variance, we have that after noise addition the backdoor inputs tend to generate similar images, while the clean inputs tend to generate diversified images. Take TrojDiff(In-D2D) attack on Cifar10 for example, $N$ is chosen as $32\times 32\times 3=3072$. Based on Lemma 3.1, we must have that $\|x_{1}-x_{2}\|_{2}\geq\|x_{3}-x_{4}\|_{2}$ under the condition of $\sigma_{In-D2D}<\frac{3071}{3072}\sigma_{c}$, where $x_{1},x_{2}\sim\mathcal{N}(x_{c},\sigma_{c}\frac{1}{\rho^{2}}I)$ and $x_{3},x_{4}\sim\mathcal{N}(x_{b},\sigma_{In-D2D}\frac{1}{\rho^{2}}I)$. However, according to Equation 14, this conditional is trivial to achieve, since $\frac{1}{\rho^{2}}\sigma_{c}\gg\frac{1}{\rho^{2}}\sigma_{In-D2D}$. This completes the proof.
Appendix H Proof of Lemma 3.2
Lemma H.1.
Let $f_{\theta}$ and $f_{\tilde{\theta}}$ be two well-trained diffusion models as defined in the Assumption G.4. Let $x^{\prime}_{T}$ follow $\mathcal{N}(0,\rho^{2}I)$. Let the clean data distribution be $q(x)\sim\mathcal{N}(x_{c},\sigma_{c}I)$. We then have:
@@
the Equation 16 is due to Assumption G.4. In particular, to obtain the generated image follows $q(x)$, the reverse process is defined as $q(x^{\prime}_{t-1}|x^{\prime}_{t})\sim\mathcal{N}(x^{\prime}_{t-1};\mu_{\tilde{\theta}}(x^{\prime}_{t},t),\Sigma_{\tilde{\theta}}(x^{\prime}_{t},t))$, where $\mu_{\tilde{\theta}}(x^{\prime}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x^{\prime}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}{\epsilon}_{t}\Big{)}$ and $\Sigma_{\tilde{\theta}}(x^{\prime}_{t},t)=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}$ (Equation 11). For the $f_{\theta}$, although the reverse process is also a gaussian distribution, $\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}(x_{t},t)\Big{)}$, $\Sigma_{\theta}(x_{t},t)=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}$ (Equation 6 and 7).
Under the Assumption G.4, $\epsilon_{\theta}$ is able to accuaratly predict the noise added on the $\sqrt{\bar{\alpha}_{t}x_{0}}$ to obtain $x_{t}$, hence the prediction of $\epsilon_{\theta}$ in Equation 17 should be $\rho\epsilon_{t}$. We have by substituting $\rho\epsilon_{t}$ into Equation 17:
By comparing Equation 19 and Equation 18 with Equation 11 , we found the mean of the reverse process is the same when inputting the $x^{\prime}_{T}$ to the $f_{\theta}$ and $f_{\tilde{\theta}}$, while the variance of $f_{\tilde{\theta}}$ is $\rho^{2}$ times larger than $f_{\theta}$. For simplicity, Let $a_{t}=\mu_{\theta}(x^{\prime}_{t},t)$ and $b_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}$, we have: $q_{\theta}(x_{t-1}|x_{t})\sim\mathcal{N}(a_{t},b_{t}I)$ and $q_{\tilde{{\theta}}}(x_{t-1}|x_{t})\sim\mathcal{N}(a_{t},b_{t}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}I)$. Hence, by the reparameterization trick, the variance of the generated $\hat{\tilde{x_{0}}}$ of $f_{\tilde{\theta}}$ is $\rho^{2}$ times greater than $f_{\theta}$. Without loss of generality, we use the Gaussian distribution to discribe the output distribution. Given the Assumption G.4, and $q(x)\sim\mathcal{N}(x_{c},\sigma_{c}I)$, $\hat{x_{0}}=f_{\theta}(x^{\prime}_{T})\sim\mathcal{N}(x_{c},\frac{\sigma_{c}}{\rho^{2}}I)$, which completes the proof. ∎
Appendix I Proof of Lemma 3.1
Lemma I.1 [10]).
Given that $x\sim\mathcal{N}(\mu,\sigma I)$, where $x$ is a $N$-dimensional vector. Then, we must have $\frac{N}{\sqrt{N+1}}\leq\sigma^{-1}\mathbb{E}(\|x\|_{2})\leq\sqrt{N}$, where $\mathbb{E}(X)$ denotes an expectation value of the random variable $X$.
Lemma I.2.
[…]
Proof.
Let $A:=\|x_{1}-x_{2}\|_{2}$, $B:=\|x_{3}-x_{4}\|_{2}$, $C:=B-A$, and $t$ be some arbitrary positive number. According to the Markov inequality, we have that
$\begin{split}Pr(C>t)\leq\frac{\mathbb{E}(C)}{t}=\frac{\mathbb{E}(B-A)}{t}=\frac{\mathbb{E}(B)-\mathbb{E}(A)}{t}=\frac{\mathbb{E}(\|x_{3}-x_{4}\|_{2})-\mathbb{E}(\|x_{1}-x_{2}\|_{2})}{t}\end{split}$ (20)
By basic calculations, it is obvious that $x_{1}-x_{2}\sim\mathcal{N}(0,\sqrt{2}\sigma_{1})$ and $x_{3}-x_{4}\sim\mathcal{N}(0,\sqrt{2}\sigma_{2})$. Then, according to Lemma I.1, we obtain that,
$Pr(C>t)\leq\sqrt{2}(\frac{\sigma_{2}\sqrt{N}}{t}-\frac{N\sigma_{1}}{t\sqrt{N+1}})\leq\sqrt{2}\frac{\sigma_{2}(N+1)-N\sigma_{1}}{t\sqrt{N+1}}$ (21)
If we have that $\sigma_{2}<\frac{N}{N+1}\sigma_{1}$, then the numerator will be always less than 0. This completes the proof. ∎

### v2

@@ Appendix I More Details about Score Distributions
$\begin{split}\Sigma_{\tilde{\theta}}(x^{\prime}_{t},t)&=1/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(\frac{\alpha_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}})=1/(\frac{\alpha_{t}-\bar{\alpha}_{t}+\beta_{t}}{\beta_{t}(1-\bar{\alpha}_{t-1})})=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}\\ \mu_{\tilde{\theta}}(x^{\prime}_{t},t)&=\frac{1}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}}(\frac{\sqrt{\alpha_{t}}}{\beta_{t}}x^{\prime}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}}{1-\bar{\alpha}_{t-1}}x_{0})/\frac{1}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}}(\frac{\alpha_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}})\\ &=(\frac{\sqrt{\alpha_{t}}}{\beta_{t}}x^{\prime}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}}{1-\bar{\alpha}_{t-1}}x_{0})\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}\\ &=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}x^{\prime}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}x_{0}\\ &=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}x^{\prime}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\frac{1}{\sqrt{\bar{\alpha}_{t}}}(x^{\prime}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}\boldsymbol{\epsilon}_{t})\\ &=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x^{\prime}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}\boldsymbol{\epsilon}_{t}\Big{)}\end{split}$
According to Lemma 1, if we add Gaussian noise to the origin input image, which results in $\mathcal{N}(0,\rho^{2}I)$, then the distribution of generated images of the diffusion model has the same mean, but a variance scaled by $\frac{1}{\rho^{2}}$, where $\rho^{2}=2$ for clean samples.
@@
Definition 12 (Backdoor Forward process).
Let $x_{0}^{b}\sim q(x_{b})$ denote a sample from target data distribution, $\delta$ denote a trigger, and $x_{T}^{b}\sim\mathcal{N}(\delta,I)$ denote the pure Gaussian noise attached by a trigger. Given the variance schedule $\{\beta_{t}\}_{t=1}^{T}$ in DDPM (Ho, Jain, and Abbeel 2020), define the forward process to diffuse $x_{0}^{b}$ to $x_{T}^{b}$ for backdoor samples:
@@
Therefore, under the Lemma 1, the distributions generated by clean and backdoor samples after noise addition are $\mathcal{N}(x_{c},\sigma^{\prime}_{c}I)$ and $\mathcal{N}(x_{b},\sigma_{b}^{\prime}I)$, respectively, where $\sigma^{\prime}_{c}=\sigma_{c}\frac{1}{\rho^{2}}$ and $\sigma_{b}^{\prime}=\sigma_{b}\frac{1}{\rho^{2}}$. In reality, the variance of clean generations $\sigma_{c}$ can be a much larger value than that of the backdoor generations. Based on the assumption that $\sigma_{c}-\sigma_{b}>\rho^{2}$, we then have that $\sigma_{c}^{\prime}>\sigma_{b}^{\prime}+1$.
This completes the proof.
Lemma 13 (Bounds on Expected Length of Gaussian Random Variable (Chandrasekaran et al. 2012)).
Given that $x\sim\mathcal{N}(\mu,\sigma I)$, where $x$ is a $N$-dimensional vector. Let $\mathbb{E}(X)$ be the expectation value of the random variable $X$. Then, we have
$\frac{N}{\sqrt{N+1}}\leq\sigma^{-1}\mathbb{E}(\|x\|_{2})\leq\sqrt{N}$
Corollary 14.
[…]
$\mathbb{E}(\|x_{1}-x_{2}\|_{2}-\|x_{3}-x_{4}\|)\geq\frac{N(\sigma_{c}-\sigma_{b})-\sigma_{b}}{\sqrt{N+1}}>0.$
@@
It is obvious that $x_{1}-x_{2}\sim\mathcal{N}(0,\sqrt{2}\sigma_{c})$ and $x_{3}-x_{4}\sim\mathcal{N}(0,\sqrt{2}\sigma_{b})$. Then, According to Lemma 13, we obtain that
$\begin{split}\mathbb{E}(\|x_{1}-x_{2}\|_{2}-\|x_{3}-x_{4}\|)&=\mathbb{E}(\|x_{1}-x_{2}\|_{2})-\mathbb{E}(\|x_{3}-x_{4}\|)\\ &\geq\sqrt{2}(\frac{N\sigma_{c}}{\sqrt{N+1}}-\sqrt{N}\sigma_{b})\\ &=\sqrt{2}\frac{N\sigma_{c}-\sqrt{N(N+1)}\sigma_{b}}{\sqrt{N+1}}\\ &\geq\sqrt{2}\frac{N\sigma_{c}-(N+1)\sigma_{b}}{\sqrt{N+1}}\\ &=\sqrt{2}\frac{N(\sigma_{c}-\sigma_{b})-\sigma_{b}}{\sqrt{N+1}}\end{split}$ (19)
Based on the conclusion from the Theorem 2, we have that $\sigma_{c}-\sigma_{b}>1$. Therefore, we obtain that,
$\mathbb{E}(\|x_{1}-x_{2}\|_{2}-\|x_{3}-x_{4}\|)>\sqrt{2}\frac{N-\sigma_{b}}{\sqrt{N+1}}\geq 0,$ (20)
where the last inequality hold because $\sigma_{b}=\mathcal{O}(N)$. This completes the proof. ∎